Obtain an equivalent system by performing the stated elementary operation on the system. Replace the third equation by the sum of itself and negative −3 times the second equation
7x+6y-15z=-14
8x-4y-6z=24
3x+10y-5z=-12

Answer by stanbon(75887)

Obtain an equivalent system by performing the stated elementary operation on the system. Replace the third equation by the sum of itself and negative −3 times the second equation
7x+6y-15z=-14
8x-4y-6z=24
3x+10y-5z=-12
-----------------------
Ans::
7x+6y-15z=-14
8x-4y-6z=24
-21x+22y+13z=-84
-----------
Cheers,
Stan H.
------------